An electric field is generated in between two parallel hundred square meter electrodes that are separated by a distance of 2 centimeters.
We have the one plate charged to 5.0 nano coulomb and the other one. Exactly equal but opposite.

An additional square shaped surface is placed in between plates
1
Expert's answer
2021-08-12T06:19:42-0400

"\Phi=ES'\cos \theta,"

"E=\frac Ud=\frac q{Cd}=\frac{qd}{d\varepsilon_0 S}=\frac q{\varepsilon_0 S},"

"\Phi=\frac{qS'}{\varepsilon_0 S}\cos \theta=8~\text{V}\cdot \text{m}."